2.001 ÷ 0.003 = __________
 2.001 ÷ 0.003 = __________
 2.001 ÷ 0.003 = __________

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

2.001 ÷ 0.003 = 667

= 2.001/0.003
= 2001/3... [divideboth numerator and denominator by 3]
= 667


Thanks